40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 354     Entry time: Wed Apr 24 13:58:51 2019
Author: Joe 
Type: Optics 
Category: Characterization 
Subject: OMC power budget and UV Epoxy Bonding of BS1 

[koji,philip,joe,liyuan,stephen]

need to add spot positions.

Mirrors: PZT11,PZT22, A14, A5

Measurement postion Power P_normalise
P_in 15.66+-0.01mV 3.251+-0.001
V_ref,lock 64+-2mV 3.22+-0.001
V_ref,unlock 2.808+-0.001 V 3.253+-0.001
P_qpd 99.5+-0.5 uW 3.24+-0.002
P_cm1 79.0+-0.5 uW 3.22+-0.002
P_cm2 76.2+-0.03 uW 3.22+-0.01
P_trans 14.55+-0.05 mW 3.22+-0.01
Vref,dark -6.286 mV +-0.01mV  

Mode matching = 97.72%

15.66-> 15.30mW coupled.

~100uW for QPD

->15.2mW in cavity

Trans = 14.55mW -> 95.7% transmission

The flat mirrors were the ones with the most scattering, so we thought about how to improve it. We tried to move the first flat mirror by pushing it with our finger so that he beam would move along the optic. We tried this a couple of times, however the second time we moved it we lost our alignment and could not retrieve it. We looked at the mirror and we could see quite a lot of newtonian rings. We could see a small fibre on the glass bread board. We cleaned the optics base and the gbb, and we could get the alignment back. The beam was aligned to the cavity, the spots no longer hit the centre of the CM2. 

We measured the power budget again.

Measurement position Power P_normalise
V_ref,lock 47mV 3.24V
P_trans 14.45+-0.005mW 3.24 +-0.003 V
V_ref,unlock 2.68+-0.001 V 3.25+-.003
     

mode matching = 1-47/2680 = 0.9824, 98.2% mode matching

same p_normalise so

15.66-> 15.34mW coupled.

~15.24mW in cavity

transmission = 14.45, so 94.8% transmission.

Koji noticed that FM1 wasn't touching the template correctly, so he re-aligned the cavity.

Afternoon session - UV Bonding (E1300201-v1 procedure 6.4.4 "Gluing" using procedure in section 7.2 "UV Gluing")

Wiped down UV PPE, UV Illuminator, and UV Power Meter

Applied Optocast 3553-LV Epoxy to sample fused silica optics, to test quantity of glue needed and to become familiar with the process and tools. Philip and Joe each created a successful bond. Joe's had 3 visible spots in the bulk of the bond. Acetone was used to scrub some residue of epoxy from the surface near the OD, which was likely cured. Short duration exposure (seconds) to acetone at the perimeter of the bond did not yield any weakening of bond.

While test pieces were bonded, Koji was making some adjustments to the cavity alignment in preparation for gluing of the steering mirror BS1.

Koji noticed that the spring clamp was causing pitch in the BS1 mirror, so he recommended that we utilize the "restrain by allen key" technique to load the mirror during curing.

Once aligned, we tried taking the BS1 mirror out of the template and then putting it back. We did this twice and both times the cavity needed realigning (with the curved mirrors as well as the input steering periscope). Why is this? Since the mirror was touching the template it should not have become misaligned right? Maybe the template moves slightly? I think before glueing in the cavity mirrors we should find out why probably? Koji took a look and claimed that a few optics may have been unconstrained. 

Planning between Koji and Joe led to placement of 5 drops of epoxy on the BS1 surface, to match the bonding area. At this point we noticed that the template was not secured very well, by poking down on it we could see it move. This might explain why we are becoming misaligned very easily. Once the prism was back on the board, Koji used allen keys to move around the prism. This was done until we could align it again (i.t looked too pitched). The beam was aligned back into the cavity, and the UV light was used to cure the bond. The reflected DC when locked was

  • pre-cured = 47mV
  • cured = 55 mV

so it looks ok still.

 

 

 

ELOG V3.1.3-